Calculate the increase in length when 1000.0cm of steel wire of linear expansitivity 1.2×10-5°c is heated from 0°c to 50°c

1
Expert's answer
2022-05-02T13:36:16-0400
"\Delta l=l_0\alpha\Delta T\\
=1000.0\:\rm cm\times 1.2\times 10^{-5}\times 50=0.60\: cm"